OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y − x2 = 11x + 30?

OpenStudy (anonymous):

You're gonna have to solve the quadratic. \[y - x^{2} = 11x + 30\] \[y = x^{2} + 11x + 30\] Substitute y for 0. \[0 = x^{2} + 11x + 30\] Can you solve the quadratic now?

OpenStudy (anonymous):

i subtract x^2 right?

OpenStudy (anonymous):

Now. You have to factor it.

OpenStudy (anonymous):

oh alright, i ended up with x= -6,-5... so its (-5,0) and (-6,0)?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

alright thanks!